The lengths of the sides of a triangle are consecutive odd integers. Find the length of the longest side if it is 20 units shorter that the perimeter.
Answer by josmiceli(19441)

Call the sides n, n+2, n+4
p = perimeter
p = 3n + 6
n + 4 is the longest side
n + 4 = 3n + 6 - 20
2n = 4 - 6 + 20
2n = 18
n = 9
n+2 = 11
n+4 = 13
The are the sides